Circular No. 6161

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

COMET C/1994 G1 (TAKAMIZAWA-LEVY) J. V. Scotti, Lunar and Planetary Laboratory, reports that observa-

tions with the 0.9-m Spacewatch telescope on Apr. 5 and 6 show a secondary nucleus of comet C/1994 G1 = 1994f, 2 mag fainter than the primary and separated by 6′′.9 in p.a. 70^{o}. No companion was noted in the previous Spacewatch images, obtained on Jan. 3. Z. Sekanina, Jet Propulsion Laboratory, finds that the observations are consistent with splitting on 1994 Sept. 4 (± 15), 105 days after perihelion, the companion’s deceleration being 21.8 (± 6.8) × 10^{−5} solar attraction. Ephemeris: 1995 Apr. 23, separation 7′′.2 in p.a. 68^{o}; May 13, 7′′.3 in 67^{o}; June 2, 7′′.2 in 65^{o}. During the Spacewatch January observation the separation would have been only 1′′.1.

γ-RAY FLARE FROM CTA 26 J. R. Mattox, University of Maryland and Universities Space Research

Association; S. Wagner, Landessternwarte Heidelberg; T. A. McGlynn, Compton Observatory Science Support Center; M. Malkan, University of California, Los Angeles; J. F. Schachter, Harvard-Smithsonian Center for Astrophysics; and P. Sreekumar, Goddard Space Flight Center, on behalf of the Compton Observatory EGRET Team, communicate: “We report the EGRET detection of quasar CTA 26 (PKS 0336–019). This opticallyviolently-variable object has z = 0.852. The flux at energy >100 MeV during Apr. 4–11 was at least an order of magnitude higher than the upper limits obtained in previous observations. EGRET is not currently observing this source, but will observe CTA 26 again during May 9–June 6. Observations at other wavelengths are of interest. Unfortunately, the sun is now only 35^{o} from CTA 26, and approaching. Contact mattox@astro.umd.edu for more information.”

COMET D/1931 R1 MPC 24933 contains observations and MPC 25030 elliptical orbital ele-

ments of a mag 15 comet discovered by C. W. Tombaugh on exposures with the 0.3-m photographic telescope at the Lowell Observatory in Sept. 1931. As with C/1931 AN [cf. MPC 24423 and 24544; previously announced by Lampland and Newman (1933, A.N. 249, 105) as a minor planet], information about D/1931 R1 became available in the late 1980s as the result of precise measurements by B. A. Skiff following D. H. Levy’s examination of the Lowell Observatory plate archive.

1995 April 13 Daniel W. E. Green